OpenStudy (anonymous):

Perform indicated operation and leave answer in the form a+bi. (7+2i)/(4-3i)

OpenStudy (anonymous):

Basically, you have to rationalize the denominator which means to get rid of 'i' in the denominator. THat means that you have to multiply by the conjugate in the denominator. \[\frac{7 + 2i}{4 - 3i} \implies \frac{7 + 2i}{4 - 3i} \times \frac{4 + 3i}{4 + 3i} \implies~?\]

OpenStudy (anonymous):

Yes...that is correct if you wanted to use a calculator...

OpenStudy (anonymous):

So it that the most simplified? Or could I simplify it down more?

OpenStudy (anonymous):

@Calcmathlete

OpenStudy (anonymous):

THat is the most simplified...I'm not surte if your teacher wants the decimal one or the fraction one. It's up to you :)

OpenStudy (anonymous):

O.k, thanks so much :)